# Is In-Context Learning Sufficient for Instruction Following in LLMs?

### v1

@@ 3 Many-shot ICL for instruction following
Table 2: Improved in-context demonstrations for Mistral-7B-v0.2-32k base model. We report the 1st-turn score on MT-Bench and the length-controlled (LC) win rates on AlpacaEval 2.0 when using different IC prompts.
@@
Urial + 17 random examples $7.22$ $8.43$
Urial + 37 random examples $7.33$ $7.18$
@@
4 Understanding ICL for instruction following
In the following, we provide further analyses of the challenges associated with ICL in the context of alignment.
4.1 Importance of question-answer matching for in-context demonstrations
Min et al. (2022) showed that using demonstrations with random labels does not significantly impair the results of ICL on classification and multiple choice tasks. We conduct a similar study for instruction following and report the results of the 1st-turn MT-Bench score in Table 3. Denoting $\{(X_{i},Y_{i})\}_{i}$ the set of in-context examples, with $X_{i}$ the query and $Y_{i}$ the corresponding ground-truth answer, we test several configurations on two sets of $\{(X_{i},Y_{i})\}_{i}$: the 3 Urial examples and the 3 Urial examples with the 3 examples found by our greedy search from Sec. 3. First, we do not use any demonstration, i.e., we directly use the original base model (Mistral-7B-v0.2-32k). As expected, this configuration performs poorly (2.93). Second, providing the questions without the answers leads to surprisingly good results ($\approx$6 MT-Bench score). Third, we permute the answers $Y_{i}$s with a circular shift of one position, so that all correct answers are still contained in the prompt but matched with the wrong question. This procedure significantly degrades the performance, especially with more examples (Urial + Greedy Search) where only a score of 1.80 is obtained (note that the minimum score is 1). Next, for each question, we sample a new answer from those provided for other instructions in the same category (in-domain). For example, a question about coding is paired with an answer from a different coding question. This approach ensures that while the content of the answer may be incorrect, its style and format remain appropriate. Although it is worse than the original one, this configuration achieves quite a good score (6.24 and 5.35). Finally, we sample answers from those of instructions from different, out-of-domain categories, so that even their style is incorrect compared to what implied by each question: this sampling leads to the worst instruction following results.
These results show that not all the conclusions from Min et al. (2022) apply to ICL used for instruction following. Most importantly, using answers with correct content and, especially, correct style is crucial for the success of ICL. This property becomes even more evident when increasing the number of examples (to 6 instead of 3). In that case, not using the correct answers significantly degrades the performance. In turn, this result suggests that ICL for instruction following works differently compared to other less open-ended tasks. This intricacy might explain the gap between ICL and fine-tuning illustrated in Sec. 3.
Table 3: Importance of question-answer matching for in-context demonstrations. We report the 1st-turn MT-Bench score of Mistral-7B-v0.2-32k when varying the structure of the in-context examples $\{(X_{i},Y_{i})\}_{i}$, where $X_{i}$ is the query and $Y_{i}$ is the corresponding ground-truth answer.
no demonstrations 2.93 2.93
$X$ only (no $Y$) 5.90 5.97
circular shift of $Y$ 5.04 1.80
in-domain random $Y$ 6.24 5.35
out-of-domain random $Y$ 4.13 1.74
correct $Y$ 6.99 7.68
4.2 Transferability of in-context examples
In Table 4, we report the performance of applying the in-context examples found by greedy search (added to Urial) on Mistral-7B-v0.2-32k (see Sec. 3) to Llama-2-7B-80k (Fu et al., 2024) and Mixtral-8x22B-v0.1-4bit (Jiang et al., 2024). Adding the new examples does not provide a consistent improvement: while it can increase the MT-Bench score (+0.47 on Llama-2-7B-80k, +0.30 on Mixtral-8x22B-v0.1-4bit), it can also, in some cases, give worse results than the original Urial. Similarly, it yields mixed results when measured by win rate on AlpacaEval 2.0, which was not optimized by the greedy search. This analysis suggests that the most effective ICL demonstrations might vary across base LLMs, potentially because of differences in their pre-training. This could further explain why Urial underperforms on multiple models in Table 1, especially on those which have become available only recently and were not used for selecting the URIAL examples.
@@ 4.2 Transferability of in-context examples
5 Conclusions and discussion
In this work, we have first illustrated that ICL via Urial provides a good baseline for instruction-following alignment, but presents a few limitations: it mostly performs slightly worse than fine-tuning on single-turn conversations and does not generalize well to multi-round ones. Focusing on single-turn evaluation, we have shown that simply adding more in-context demonstrations is not sufficient to reliably improve performance, unlike what happens for less open-ended tasks (Agarwal et al., 2024; Bertsch et al., 2024). Even with more sophisticated approaches, fully closing the gap between ICL and IFT does not appear easy. We have provided several analyses of why this could be the case, showing that ICL for instruction following presents different behavior compared to the tasks discussed in Min et al. (2022): in our case, the demonstrations need to be carefully chosen and of high quality, with correct answers to each question. We conjecture that via ICL, the LLM can learn to infer the response style, but the overall learning signal is not sufficiently strong to benefit from a large amount of examples, despite the very long context windows which have become available only recently. Interestingly, a few incorrect ICL demonstrations are sufficient to severely degrade the performance.
In general, LLMs still require many different post-training steps beyond simple instruction following: domain-specific continued pre-training, long-context fine-tuning, fine-tuning for function calling, safety training against jailbreaks, prevention of prompt injections, etc. The difficulty of achieving even the baseline instruction-following alignment through ICL raises questions about the feasibility of obtaining more refined capabilities without fine-tuning.
Acknowledgements
We are grateful to OpenAI for providing us API credits and access to GPT-4-Base as part of the OpenAI Researcher Access Program. M.A. was supported by the Google Fellowship and Open Phil AI Fellowship.
Appendix A Experimental Details
A.1 Datasets and base models
We select high-quality demonstrations for the base models to learn at test time from open-sourced instruction fine-tuning datasets: (a) Evol-Instruct-70k (Xu et al., 2024) contains 70k training examples with varying complexity and is well-known for the use to build the series of WizardLM models, and (b) OpenHermes-2.5 (Teknium, 2023) is a continuation of the OpenHermes-1 dataset, at a much larger scale, much more diverse, and much higher quality compilation. We employ Evol-Instruct-70k dataset, which leads to a better open-sourced aligned model in terms of the MT-Bench score, for the Random scaling experiment and use OpenHermes-2.5 for the Diversity scaling experiment since it contains annotations for the data source.
  1. 1.
    Mistral-7B-v0.2 (Jiang et al., 2023) is the newest 7B model from Mistral when we are writing and its instruct version significantly outperforms Mistral-7B-Instruct-v0.1 on MT-Bench. The trained context length of Mistral-7B-v0.2 is 32k tokens.
  2. 2.
  3. 3.
We use the same inference configuration as Urial (Lin et al., 2024). Concretely, we employ greedy decoding for all models, including base and instruction fine-tuned models, to maximize reproducibility and secure a fair and robust evaluation. Moreover, we use a repetition penalty on base models to prevent degeneration, 1.0 for MoE models and 1.15 for the remaining models.
A.2 Evaluation
MT-Bench (Zheng et al., 2023), consists of 80 high-quality and challenging questions that have multi-round interaction, designed to examine the multi-turn conversation and instruction-following capability of models. It features 8 common categories of user prompts: coding, math, reasoning, extraction, roleplay, writing, humanities/social science, and STEM.
AlpacaEval 2.0 (Li et al., 2023) provides 805 test instructions, on which we generate new responses using the target model, and then calculate the score by competing with the baseline model (i.e., GPT-4-Turbo) judged by a designated automatic evaluator. We apply the AlpacaEval 2.0 benchmark in our experiments to ensure that the effective demonstrations found through greedy search don’t overfit to MT-Bench.
A.3 Greedy search
We start with constructing a dataset that solely consists of high-quality examples from Evol-Instruct-70k. We then sample 100 high-quality examples as the initial search space. First, we employ GPT-3.5-Turbo as the judge to evaluate the quality of each data point in Evol-Instruct-70k with a prompt template as shown in Fig. 3. A thorough examination of the Evol-Instruct-70k dataset reveals a score distribution (Fig. 4), in which over 14k examples have a score of at least 4.5. Subsequently, we truncate the entire dataset by setting up a threshold of 4.5, resulting in a smaller dataset that is exclusively comprised of high-quality examples as indicated by GPT-3.5-Turbo.
Figure 3: The prompt template for evaluating the quality of data points using GPT-3.5-Turbo as the judge.
Next, we randomly sample 100 examples and then create 100 new prompts by adding each one of the sampled examples, as the fourth demonstration, to the original 3-shot Urial prompt. We evaluate the resulting instruction-following performance for each prompt using MT-Bench but with GPT-4-Turbo as the judge, and select the best fourth example based on the 1st-turn score. We continue our greedy search for the proper fifth and sixth demonstration in a restricted search space. Due to computational cost, we keep reducing the search space via applying a threshold, which is decided based on the 1st-turn MT-Bench score distribution we get in the previous round of searching.
A.4 Scaling experiments
The recently released LLMs with increasingly large context windows allow using many more shots for ICL than are typically used. Prior works (Agarwal et al., 2024; Bertsch et al., 2024) have shown that many-shot ICL, compared to few-shot ICL, can make task-specific fine-tuning less essential and allows LLMs to tackle a wider range of tasks without specification. Therefore, we want to examine if the instruction-following performance of base LLMs could benefit from many-shot ICL through various scaling experiments. In the following, we add more details about these scaling experiments, including the prompt template and data selection for ICL.
Figure 4: The distribution of data quality for Evol-Instruct-70k dataset as measured by using GPT-3.5-Turbo as the judge.
Random.
We construct the set of in-context demonstrations based on the high-quality dataset we create from Evol-Instruct-70k. Through randomly sampling from the high-quality dataset multiple times, we generate a series of in-context demonstration sets that each contains $N$ examples, where $N\in\{2,7,17,27,37,47\}$, and insert the $N$ demonstrations into the prompt template for ICL as shown in Fig 5. Note that the prompt template is strictly following the one used in Urial paper and we only replace the demonstrations used for the ICL purpose.
Greedy.
One could perceive the greedy search results of the proper fourth example as an alternative method of evaluating the data quality. An example adding to the context as the fourth demonstration results in a higher 1st-turn MT-Bench score, indicating its high quality in our ICL scenario. In this way, we can sort the 100 high-quality examples in descending order based on the score as evaluated by GPT-4-Turbo. Thus it allows us to create another series of in-context demonstration sets by selecting top-$N$ examples for each $N\in\{2,7,17,27,37,47\}$. The resulting mixture of in-context demonstrations is then placed in the corresponding location of the prompt template (Fig. 5).
Diversity.
Agarwal et al. (2024) suggest including diverse ICL examples could potentially lead to better generation, a stronger capability to handle more complex instructions and greater control over model outputs. Our work revisits this finding in the instruction-following task by scaling up examples from different sources. Concretely, we employ stratified sampling over the source annotation of OpenHermes-2.5 to collect $N$ diverse ICL examples, where $N\in\{3,5,10,20,30,40,50\}$. We get rid of the Urial demonstrations in this experiment since their source is unknown.
Figure 6: The prompt template for doing Unsupervised ICL in our work. Note that the words in orange are solely for illustration purposes and do not appear in the real prompts. The mixture of new demonstrations we test in our experiments is inserted before Urial demonstrations.
Unsupervised.
Motivated by encouraged results with unsupervised ICL in Agarwal et al. (2024), we examine the efficacy of scaling up the number of unsupervised queries. For simplicity, we reuse instructions from the experiment of scaling up diverse examples and remove corresponding answers. Specifically, the prompt template for unsupervised ICL scaling experiment consists of: (a) general rules that the model should follow when generating answers, (b) multiple instructions without answers, and (c) a few-shot prompt with ground truth answers for the desired output format and the Urial prompt is used in our work. We show the exact prompt template in Fig. 6.
Appendix B Additional experiments

### v2

@@ 3.1 Scaling up the number of in-context demonstrations
Table 3: Improved in-context (IC) alignment by adding to Urial prompt the demonstrations found via greedy search. We find the optimal IC prompts in an incremental way by selecting query-response pairs from the high-quality SkillMix dataset. We report the 1st-turn score on MT-Bench and the length-controlled (LC) win rates on AlpacaEval 2.0 when using different IC prompts.
Model Mistral-7B-v0.2 Llama-3.1-8B
MT-Bench (1st) AlpacaEval 2.0 MT-Bench (1st) AlpacaEval 2.0
Urial (3 examples) $7.00$ $8.22$ $6.95$ $7.28$
Urial + greedy search (1 ex.) 7.52 $7.53$ $7.61$ 8.61
Urial + greedy search (2 ex.) $7.47$ $7.78$ $7.77$ $8.16$
Urial + greedy search (3 ex.) $7.43$ 8.55 7.81 $8.19$
3.2 Greedy search for effective demonstrations
Given the limited success of adding random demonstrations to Urial, we try to greedily maximize the MT-Bench score by testing 100 high quality instructions from SkillMix as the 4th additional example to Urial. For each resulting ICL prompt, we compute the MT-Bench score with GPT-4-Turbo as judge instead of GPT-4 (used in MT-Bench) as the former is faster, cheaper, and helps mitigate potential overfitting to the benchmark score. We then repeat this procedure sequentially to find a 5th and a 6th demonstration, restricting the search space at each round to only instructions leading to a high enough MT-Bench score to reduce the computational cost (see details in App. A.3).
We add the results of the greedy search to the plots in Fig. 3, computing the true MT-Bench (i.e., with GPT-4 as judge). For both base models, the 4th example found by greedy search is sufficient to match the best 1st-turn score achieved by scaling the in-context examples with random demonstrations (see Sec. 3.1). For Llama-3.1-8B, the 5th and 6th demonstrations further improve the score, while they are not helpful for Mistral-7B-v0.2. In Table 3, we provide the details of these results: the 4th ICL example yields a significant improvement over Urial, e.g., from 6.95 to 7.61 for Llama-3.1-8B, with only a further +0.20 given by the 5th and 6th. Overall, this evaluation further indicates that the improvement in instruction following one can achieve with in-context alignment quickly saturates when increasing the number of examples. In Fig. 4, we show the distribution of the scores (with GPT-4-Turbo as judge) at step of the greedy search when using Llama-3.1-8B as base model (the analog for Mistral-7B-v0.2 in App. C.1). Most demonstrations improve the score when added as the 4th example on top of Urial (as shown in the first plot), but no further progress is observed with additional demonstrations.
We further test the IC prompts found by the greedy search on the AlpacaEval 2.0 (Li et al., 2023a) benchmark, where we measure length-controlled win rate. As shown in Table 3, using the examples given by greedy search leads to better results than plain Urial on AlpacaEval 2.0, without directly optimizing for it. This improvement indicates that the found IC prompt does not completely overfit to the MT-Bench score, although the gains are less consistent than on MT-Bench. For Mistral-7B-v0.2, Urial with 3 greedy search examples is the only configuration that outperforms Urial, while for Llama-3.1-8B just one additional demonstration yields the best win rate, with a significant improvement over Urial (7.28 vs 8.61).
4 A Comparison of ICL vs IFT for Instruction Following
(a) Mistral-7B-v0.2
(b) Llama-3.1-8B
Figure 5: Comparison of ICL vs IFT for alignment in the low data regime. We measure the alignment performance of different settings for Mistral-7B-v0.2 and Llama-3.1-8B using the MT-Bench score. IFT with more demonstrations keeps improving the alignment performance, almost bridging the gap between the base model and its aligned counterpart. IFT-aligned models perform well on multi-turn conversations, unlike with ICL. Finally, data quality has significant impact on both IFT and ICL: the higher-quality SkillMix leads to better performance than Evol-Instruct.
The results from the previous sections strongly suggest that in-context alignment might not be as competitive as sophisticated, computationally heavy alignment techniques. However, it is unclear how it compares to more lightweight approaches, such as IFT, when applied to a small number of examples. In the low-data regime, the choice between ICL and IFT can be viewed as a decision between allocating resources for fine-tuning (which permanently modifies the model’s weights) or increasing inference time by adding the in-context prompt each time.
Setup. We test IFT on the same number of instructions from SkillMix as used in the many-shot ICL experiments in Sec. 3, ranging from 3 to 250 (the maximum allowed by the context length of the base models). Moreover, we scale the dataset up to 4k, that is the entire SkillMix. Each training run is repeated over multiple seeds, resulting in different training datasets (more details in App. A.4).
Results on Mistral-7B-v0.2. The top part of Fig. 5 shows that IFT and ICL on SkillMix (violet and green curves respectively) perform almost identically in 1st-turn score until 50 examples are used. Beyond this point, increasing the number of training examples consistently improves the performance of IFT. The strong performance of IFT in the low-data regime is particularly surprising, as one might expect overfitting when training on very few examples (as few as 3) over several epochs, yet this does not occur. Finally, ICL and IFT show nearly opposite trends for the 2nd-turn score: increasing the dataset size benefits IFT (almost reaching the performance of the instruct model) but detrimental to ICL. This behavior suggests that, while IFT is less flexible due to the model weights being updated, it generalizes better to tasks different from those used for alignment (recall that the instructions in SkillMix are single-turn only).
Results on Llama-3.1-7B. As shown in the bottom part of Fig. 5, ICL outperforms IFT when using between 3 and 30 demonstrations, though IFT remains effective even with these few examples on this base model. However, we note that IFT matches or exceeds the best performance of ICL (obtained with 20-30 examples) only when using 2k or 4k examples (the entire SkillMix), which represents two orders of magnitude more data. This result confirms that ICL with high-quality data is a viable alternative to IFT when only a limited number of demonstrations are available. Finally, the observations for the 2nd-turn scores are consistent with those for Mistral-7B-v0.2, with IFT consistently outperforming ICL.
Effect of data quality. Finally, we test the effect of using lower-quality instructions compared to SkillMix. In this experiment, we repeat the ICL vs. IFT comparison using random demonstrations from Evol-Instruct-70k (Xu et al., 2024), and show the results in Fig. 5. With Evol-Instruct data, both ICL (red curve) and IFT (yellow curve) perform significantly worse than their counterparts using SkillMix. This trend is consistent across the number of examples, base models, and for both single-turn and multi-turn instructions. Interestingly, the performance gap between IFT on SkillMix and Evol-Instruct is significantly smaller on Llama-3.1-8B than on Mistral-7B-v0.2, suggesting that better pre-trained models may partially compensate for lower-quality fine-tuning data. Finally, we notice that ICL with 3 examples from Evol-Instruct gets worse 1st-turn scores than Urial (with also 3 examples), whereas ICL with 3 examples from SkillMix matches (on Mistral-7B-v0.2) or outperforms (on Llama-3.1-8B) Urial (see also discussion in Sec. 3.1). These observations further confirm the importance of instruction quality for alignment, whether in the context of in-context learning or instruction fine-tuning.
5 Conclusions
In this work, we have first illustrated that ICL via Urial is a good baseline for instruction-following alignment, but with a few limitations: it typically performs slightly worse than IFT on single-turn conversations and does not generalize well to multi-round ones. Next, we have uncovered the key components for IC alignment: e.g. the decoding parameters alone can, surprisingly, make base models reasonably good at instruction following. Adding in-context high-quality demonstrations improves performance beyond what previously suggested (Lin et al., 2024), but not enough to reach LLMs aligned with sophisticated methods (e.g., RLHF). We conjecture that via ICL, the LLM can learn to infer the response style, but the overall learning signal is not sufficiently strong to benefit from a large amount of examples, despite the long context windows of recent LLMs.
Moreover, to the best of our knowledge, we have provided the first systematic comparison of ICL and IFT for instruction following when using the same (small) number of demonstrations. Surprisingly, the two approaches are roughly equivalent in terms of single-turn conversation performance. However, models trained via IFT, unlike ICL, can generalize to multi-round conversations, which are out-of-distribution compared to the examples used for alignment. Overall, our work provides a deeper and more complete understanding of how in-context alignment works, as well as of its potential and limitations, in particular in comparison to fine-tuning. This comprehension might be the basis for future work aimed at using ICL to efficiently customize LLMs without the need of fine-tuning, as well as exploring the fundamental differences between base and aligned models.
Limitations. We perform greedy search with limited computational resources, e.g. the set of candidate examples is relatively small. We expect that more effective ICL prompts could be identified with additional resources. Moreover, our focus is on instruction-following tasks, but other types of alignment, such as safety-oriented tasks, may also be of interest.
Acknowledgements
We are grateful to OpenAI for providing us API credits and access to GPT-4-Base as part of the OpenAI Researcher Access Program. M.A. was supported by the Google Fellowship and Open Phil AI Fellowship. This work was supported by the Swiss National Science Foundation (grant number 212111) and by an unrestricted gift from Google.
Appendix A Experimental Details
A.1 Datasets and base models
We select instruction-following demonstrations for the base LLMs to learn in context from open-sourced instruction fine-tuning datasets: (a) SkillMix-4k (Kaur et al., 2024) is an automated approach for synthesizing diverse, high-quality IFT data. It primarily involves two stages: first leveraging the LLMs to propose a set of critical ”skills” for instruction-following, from which a pair of skills are randomly chosen to facilitate synthetic data generation based on powerful LLMs. (b) Evol-Instruct-70k (Xu et al., 2024) contains 70k training examples with varying complexity and is well-known for the use to build the series of WizardLM models. SkillMix-4k is the primary dataset for all experiments in the paper and Evol-Instruct-70k is solely used for scaling experiments and ablation study of data quality.
  1. 1.
    Mistral-7B-v0.2 (Jiang et al., 2023) has 7.3 billion model parameters. On many established benchmarks, it outperforms LLMs that have significantly more parameters, such as Llama-2-13B and Llama-1-34B. The trained context length of Mistral-7B-v0.2 is 32k tokens.
  2. 2.
    Llama-3.1-8B (Dubey et al., 2024) is the newest and most powerful 8B model from Meta when we are writing, and it supports multilingual dialogue use cases. It supports input texts containing up to 128k tokens.
In some experiments, the following base LLMs are used:
  1. 1.
  2. 2.
[…]
A.2 Evaluation
MT-Bench (Zheng et al., 2023), consists of 80 high-quality and challenging questions that have two-round interaction, designed to examine the multi-turn conversation and instruction-following capability of models. It features 8 common categories of user prompts: coding, math, reasoning, extraction, roleplay, writing, humanities/social science, and STEM.
AlpacaEval 2.0 (Li et al., 2023a) provides 805 test instructions, on which we generate new responses using the target model, and then calculate the score by competing with the baseline model (i.e., GPT-4-Turbo) judged by a designated automatic evaluator. We apply the AlpacaEval 2.0 benchmark in our experiments to ensure that the effective demonstrations found through greedy search don’t overfit to MT-Bench.
A.3 Greedy search
Firstly, we randomly sample 100 examples from the high-quality IFT dataset, SkillMix-4k, and then create 100 new 4-shot prompts by adding each one of the 100 sampled examples, as the fourth demonstration, to the original 3-shot Urial prompt. We evaluate the resulting instruction-following performance for each prompt on MT-Bench but with GPT-4-Turbo as the judge since GPT-4-Turbo is cheaper and faster than GPT-4 (the original LLM judge for MT-Bench evaluation), and also has a high correlation with human judgment. Finally, the example with the highest 1st-turn MT-Bench score is chosen.
We continue our greedy search for the proper fifth and sixth demonstration in a more restricted search space due to the heavy computational cost of the search process. We keep reducing the search space by applying a threshold of 1st-turn MT-Bench score (i.e., 6.2 in our experiments). Similarly, we add each example from the new search space on top of the best (N-1)-shot prompt found in previous steps and generate a set of new N-shot prompts. Then we run the MT-bench evaluation with GPT-4-Turbo as the judge for multiple times and select the best (N+1)-shot prompt.
A.4 Scaling experiments
The recently released LLMs with increasingly large context windows allow using many more shots for ICL than are typically used. Prior works Agarwal et al. (2024); Bertsch et al. (2024) have shown that many-shot ICL, compared to few-shot ICL, can make task-specific fine-tuning less essential and allows LLMs to tackle a wider range of tasks without specification. Therefore, we want to examine if the instruction-following performance of base LLMs could benefit from many-shot ICL through various scaling experiments, and systematically compare ICL with IFT. In the following, we add more details about these scaling experiments, including the prompt template, data selection for both ICL and IFT, and training hyper-parameters for IFT experiments.
ICL Random.
We construct the set of in-context demonstrations based on the high-quality data we select from SkillMix-4k or Evol-Instruct-70k dataset. Through randomly sampling from the high-quality dataset for multiple times, we generate a series of in-context demonstration sets that each contains $N$ examples, where $N\in\{0,7,17,27,37,47\}$ for Mistral-7B-v0.2 model (32k context length) and $N\in\{0,7,17,27,37,47,97,147,197,247\}$ for Llama-3.1-8B model (128k context length), and insert the $N$ demonstrations into the prompt template for ICL as shown in Fig 6. Note that the prompt template is strictly following the one used in Urial paper and we only replace the demonstrations used for the ICL purpose. The average performance and standard deviations are computed over 5 random seeds.
ICL Random (no URIAL).
We generate a series of in-context demonstration sets by randomly sampling from the high-quality IFT dataset, and each set contains $N$ examples, where $N\in\{3,10,20,30,40,50\}$ for Mistral-7B-v0.2 model (32k context length) and $N\in\{3,10,20,30,40,50,100,150,200,250\}$ for Llama-3.1-8B model (128k context length). URIAL examples will not be added to the context, so it ensures the total number of in-context examples are the same as the Random group. The average performance and standard deviations are computed over 5 random seeds.
ICL Greedy Search.
Following the procedure we mention in Sec. A.3, ideally we can get as many optimal examples as we want if we have sufficient OpenAI API credits. Thus it allows us to create another series of in-context demonstration sets by selecting another $N$ examples for each $N\in\{1,2,3\}$ in the paper. The resulting mixture of in-context demonstrations is then placed in the corresponding location of the prompt template as shown in Fig. 6.
Table 4: Details of training hyperparameters for IFT experiments.
Data Size # GPUs Epochs LR LR Scheduler Batch Size Context Win. Len. WD Warmup Rate
Mistral-7B-v0.2
3 2 6 2e-6 Cosine 2 2048 0.01 0.03
10 4 6 2e-6 Cosine 8 2048 0.01 0.03
20 4 6 2e-6 Cosine 8 2048 0.01 0.03
30 4 6 2e-6 Cosine 8 2048 0.01 0.03
40 4 6 2e-6 Cosine 8 2048 0.01 0.03
50 4 6 2e-6 Cosine 8 2048 0.01 0.03
100 4 6 2e-6 Cosine 8 2048 0.01 0.03
300 4 6 2e-6 Cosine 8 2048 0.01 0.03
1000 4 15 2e-6 Cosine 128 2048 0.01 0.03
2000 4 15 2e-6 Cosine 128 2048 0.01 0.03
4000 4 15 2e-6 Cosine 128 2048 0.01 0.03
Llama-3.1-8B
3 2 6 4e-6 Cosine 2 2048 0.01 0.03
10 4 6 4e-6 Cosine 8 2048 0.01 0.03
20 4 6 4e-6 Cosine 8 2048 0.01 0.03
30 4 6 4e-6 Cosine 8 2048 0.01 0.03
40 4 6 4e-6 Cosine 8 2048 0.01 0.03
50 4 6 4e-6 Cosine 8 2048 0.01 0.03
100 4 6 4e-6 Cosine 8 2048 0.01 0.03
300 4 6 4e-6 Cosine 8 2048 0.01 0.03
1000 4 15 4e-6 Cosine 128 2048 0.01 0.03
2000 4 15 4e-6 Cosine 128 2048 0.01 0.03
4000 4 15 4e-6 Cosine 128 2048 0.01 0.03
IFT.
We run instruction fine-tuning the base LLMs (Mistral-7B-v0.2 and Llama-3.1-8B) either with tens of examples (i.e., few-sample regime) or thousands of examples. The training examples are randomly sampled from existing IFT datasets, SkillMix-4k and Evol-Instruct-70k. The mean score and standard deviations are calculated over multiple random seeds. More specifically, we use 5 random seeds for Mistral-7B-v0.2 model and 3 random seeds for Llama-3.1-8B model due to a restricted compute budget. In particular, since the size of SkillMix dataset is 4k, the randomness of IFT with 4k examples primarily comes from optimization process and scoring evaluation with GPT-4, otherwise part of randomness also comes from data sampling process. We list the details of training hyper-parameters used in IFT experiments in Table. 4. We always select the last model checkpoint to run evaluation.
Table 5: Improved in-context demonstrations selected from Evol-Instruct-70k dataset for Mistral-7B-v0.2 base model. We report the 1st-turn score on MT-Bench and the length-controlled (LC) win rates on AlpacaEval 2.0 when using different IC prompts.
@@ A.4 Scaling experiments
Appendix B Transferability of In-Context Prompts
In Table 6, we report the performance of applying the in-context examples found by greedy search (added to Urial) on Mistral-7B-v0.2 and Evol-Instruct-70k dataset (see Table. 5) to Llama-2-7B-80k (Fu et al., 2024) and Mixtral-8x22B-v0.1-4bit (Jiang et al., 2024). Adding the new examples does not provide a consistent improvement: while it can increase the MT-Bench score (+0.47 on Llama-2-7B-80k, +0.30 on Mixtral-8x22B-v0.1-4bit), it can also, in some cases, give worse results than the original Urial. Similarly, it yields mixed results when measured by win rate on AlpacaEval 2.0, which was not optimized by the greedy search. This analysis suggests that the most effective ICL demonstrations may vary across base LLMs, potentially because of differences in their pre-training. This could further explain why Urial underperforms on multiple models in Table 1, especially on those which have become available only recently and were not used for selecting the URIAL examples.
@@ Appendix B Transferability of In-Context Prompts
(a) Search for 4th example
(b) Search for 5th example
(c) Search for 6th example
Figure 7: The distribution of the 1st-turn MT-Bench score (GPT-4-Turbo as judge) on Mistral-7B-v0.2 obtained by adding multiple instructions from SkillMix (Kaur et al., 2024) as a 4th (a); 5th (b); 6th (c) demonstration to Urial. The dashed lines of various colors refer to the 1st-turn MT-Bench score of the obtained searching results. A majority of the 4th examples have positive contributions to the model’s instruction-following performance, but the improvement quickly diminishes when running the greedy search for more optimal demonstrations.
Appendix C Additional experiments
C.1 Greedy Search on Mistral-7B-v0.2
In this section, we show the distribution of the resulting scores (with GPT-4-Turbo as the judge) for each step of the greedy search on SkillMix-4k dataset when using Mistral-7B-v0.2 as the base model. Similar to the finding on Llama-3.1-8B model (see Sec. 3.2), the majority of demonstrations increase the alignment performance, as measured by 1st-turn MT-Bench score, when added as the fourth example on top of URIAL, and few progress is made with more demonstrations. We show more details of greedy search results on Mistral-7B-v0.2 in Fig. 7.
C.2 More experiments on decoding schemes
[…]
(a) No ICL / $\text{rp}=1.0$
(b) ICL with URIAL / $\text{rp}=1.0$
(c) Instruct model / $\text{rp}=1.0$
(d) No ICL / $\text{rp}=1.15$
(e) ICL with URIAL / $\text{rp}=1.15$
(f) Instruct model / $\text{rp}=1.15$
Figure 8: The 1st-turn MT-Bench scores of model generations with and without ICL across different decoding schemes. We mainly consider two hyper-parameters: temperature and top-p. The heatmaps show the answering quality of the base model with and without URIAL in the context is sensitive to the setups of decoding schemes. Surprisingly, with proper decoding parameters, the base model alone is already capable of following instructions.
[…]
Figure 9: The 1st-turn MT-Bench scores of Llama-3.1-8B generations with and without URIAL in the context across different decoding schemes. We mainly consider two hyper-parameters: temperature and top-p.
C.3 A Breakdown examination of URIAL-aligned models